Following is the current balance sheet for a local partnership of doctors:

  

  Cash and current assets $ 59,000      Liabilities $ 54,000   
  Land 159,000      A, capital 34,000   
  Building and equipment (net) 162,000      B, capital 54,000   
  C, capital 104,000   
  D, capital 134,000   
     Totals $ 380,000         Totals $ 380,000   

  

The following questions represent independent situations:

a.

E is going to invest enough money in this partnership to receive a 20 percent interest. No goodwill or bonus is to be recorded. How much should E invest?

      

b.

E contributes $90,000 in cash to the business to receive a 20 percent interest in the partnership. Goodwill is to be recorded. Profits and losses have previously been split according to the following percentages: A, 30 percent; B, 10 percent; C, 40 percent; and D, 20 percent. After E makes this investment, what are the individual capital balances?

      

c.

E contributes $52,000 in cash to the business to receive a 20 percent interest in the partnership. Goodwill is to be recorded. The four original partners share all profits and losses equally. After E makes this investment, what are the individual capital balances?

      

d.

E contributes $52,000 in cash to the business to receive a 18 percent interest in the partnership. No goodwill or other asset revaluation is to be recorded. Profits and losses have previously been split according to the following percentages: A, 10 percent; B, 30 percent; C, 20 percent; and D, 40 percent. After E makes this investment, what are the individual capital balances?

      

e.

C retires from the partnership and, as per the original partnership agreement, is to receive cash equal to 115 percent of her final capital balance. No goodwill or other asset revaluation is to be recognized. All partners share profits and losses equally. After the withdrawal, what are the individual capital balances of the remaining partners?
Answer #1

A) How much should E invest

E have to invest 20% of the partnership balances

orginal capital

capital of A 34000

B 54000

C 104000

D 134000

Total capital 326000

Capital if E invest =326000 + E 's investment

Then,

E's investment = 20%(326000+Ei)

= 65200+0.20 Ei

Ei-0.20 Ei = 65200

Ei = 65200/0.80

E's investment = 81500

Thus investment of E is 81500

B)what are the individual capital balances

E's contribution = 90000

E's interest in the partnership = 20%

thus the implied value of the partnbership = 450000 (that is, 90000*100/20)

Then the capital balance of the firm after E's contribution

Earlier total capital = 326000

+ E's contribution = 90000

Total capital = 416000

Goodwill

Implied value- capital after investment

= 450000-416000

=34000

Allocation of the Goodwill among A,B,C,D at the ratio of 30:10:40:20

A= 34000*30/100 = 10200

B= 34000*10/100 = 3400

C= 34000*40/100 = 13600

D= 34000*20/100 = 6800

Revised capital balance

A B C D E
Orginal Balance 34000 54000 104000 134000 -
Good will 10200 3400 13600 6800 -
Investment - - - - 90000
Current Revised Capital 44200 88000 117600 140800 90000

C) E contributes = 52000

E's interest in the partnership   = 20%

then the implied value of the partnership = 52000*100/20 = 260000

Then the capital balance of the firm after E's contribution

Earlier total capital = 326000

+ E's contribution = 52000

Total capital   = 378000

Here the implied value of the firm is less than the total capital then E has also contributed as good will

E's investment = 20%(326000+Ei)

= 65200+0.20 Ei

Ei-0.20 Ei = 65200

Ei = 65200/0.80

E's investment = 81500

E's contributionforn the partnership = 81500

E contributed as cash = (52000)

then goodwill contributed by E = 29500

Thus E's Capital = 81500

E' s interest in the partnership = 20%

Thus the implied value of the partnership = 81500*100/20 =407500

Capital balance after E's investment = 326000 + 81500

= 407500

Here the old partners are not getting any amount of goodwills and thus it is not be shared among them, thus the revised capital of the partnership is

A B C D E
Orginal Balance 34000 54000 104000 134000
Investment 81500
Capital Balance 34000 54000 104000 134000 81500

D)

E contributes = 52000

E's interest in the partnership = 18%

implied value of the partnership = 52000*100/18 = 288888.8

Capital after E's investment = 326000+52000

= 378000

E's capital balance = 378000*18/100 = 68040

Less E's contribution = (52000)

Bonus Given to E = 16040

Bonus Born by other partners by A,B,C,D as 10:30:20:40

A = 16040*10% = 1604

B = 16040*30% = 4812

C = 16040*20% = 3208

D = 16040*40% = 6416

Revised capital

A B C D E
Orginal capital 34000 54000 104000 134000 -
Investment - - - - 52000
Bonus (1604) (4812) (3208) (6416) -
Capital Balances 32396 49188 100792 127784 52000

As here there is no goodwill in the contribution of E , thus the difference is considered as the bonus contributed by the existing shareholders. The bonus amount will be deducted from their capitals

E) While C retires

C's capital orginal balance = 104000

Amount to be collected by C = 104000*115% = 119600

Bonus given to C = 119600-104000

= 15600

The bonus amount to pay to the C is born by the other three partners A,B,D equally that is, 15600/3 = 5200

Revised capital balance of the partnership

A B C D
Orginal balance 34000 54000 104000 134000
Bonus (5200) (5200) 15600 (5200)
Payments - - 119600 -
Capital Balances 28800 48800 - 128800